Grade 12th passElectric Current

L= 3mH
i= 80cos150t
find the terminal voltage and the energy stored in it

himanshi
7 Years agoGrade 12th pass

1 Answer

Sanjit
7 Years ago
V= L.dI/dt
    = 3 x 10-3 (80 x 150 x -sin 150t)
    = 3600 x 10-3 x -sin 150t
    = -3.6 sin 150t
    
For energy stored......
    U=(1/2) LI2
        =(1/2) (3 x 10-3) (80 cos 150t)2
       =(1/2)(3x10-3) (80 x80 x cos2 150t)
       =(1/2)(3 x 10-3) (6400 cos2 150t)
   U =9.6 cos150t J